OpenStudy (anonymous):

You need a 20% alcohol solution. On hand, you have a 400 mL of a 15% alcohol mixture. How much pure alcohol will you need to add to obtain the desired solution?

OpenStudy (anonymous):

15% is basically the following fraction:\[15\% = \frac{.15}{1}=\frac{.15}{1}\cdot\frac{400}{400} = \frac{60}{400}\]Now, add \(x\) amount of milliliters of alcohol (this will increase BOTH the numerator (amount of alcohol) and denominator(amount of liquid)).\[\frac{60+x}{400+x}\]That's a start for you.

OpenStudy (anonymous):

I'm still really confused. When we went over these types of problems in class I remember cross multiplying?